There is about an 66.6% chance that a new drug will cure AIDS. Suppose that the drug is tested with four patients. Find the approximate probability that one or two of the four patients will be cured by the drug.

0.45 X Bin(4,2/3) 0.4 0.1 0.05 0 1
Answer #1

P(X = 1 or 2) = P(X = 1) + P(X = 2) = 0.1 + 0.29 = 0.39